# CubeDAgger: Improved Robustness of Interactive Imitation Learning without Violation of Dynamic Stability

### v1

@@ IV Simulations
In order to statistically verify the performance of the proposed method, CubeDAgger, the following simulations are conducted. The important criteria are threefold: i) the difference between the expert’s action $a$ and the executed action $a^{c}$ during the data collection process; ii) the degree of degradation of control performance during the data collection process; iii) the robustness of the learned policies.
IV-A Tasks
Three tasks with different robots simulated in Mujoco are performed. The first Pusher, where a manipulator pushes an object to its destination without grasping it, might cause the object to slide into hard-to-push positions by wrong actions. The second HalfCheetah, where a 2D legged robot moves forward while standing on its head [^2] This behavior was achieved by changing the weights of rewards, is unstable and sensitive to mistakes. The third HexaAnt, which has a multi-degree-of-freedom system with three joints in each of its six legs for walking foward, is easy to intervene in state transitions and stability with various action patterns. These tasks were trained multiple times with the RL algorithm [30], and the best policies are selected as their experts. Since the experts maximized the sum of rewards, the control performance is evaluated by it.
IV-B Models
The policy of the imitation agent is approximated by neural networks with two fully-connected layers (each with 100 neurons) in the hidden layer. To make it a lightweight ensemble model, only the output layer is separated to take $K=10$ outputs (i.e. the mean and log scale of Gaussian distribution), while sharing the input and hidden layers [31]. This policy ensembles (and other learnable parameters) are optimized using AdaTerm [24], a stochastic gradient descent method that is robust to noise and outliers. The dataset $D$ is assumed to have the sufficient capacity to hold all the data during the experiments, and all the data in $D$ are uniformly randomly replayed once at the end of each episode with batch size $B=50$ each.
IV-C Comparisons
@@
Here, for fair comparison, the same threshold value of $\bar{\sigma}=0.1$ is used for C1, C2, and C3.
IV-D Results
The difference $|a-a^{c}|$ (labeled Diff) and the sum of rewards (labeled Score) at the time of data collection are plotted in Fig. 3. Apparently, EV1 frequently performed actions different from those of the experts, causing large drops in the scores. This behavior would involve significant risk during data collection in real-world robot applications, even if it is needed to learn how to recover from (accumulated) errors in case of failure. On the other hand, EV2 mostly prioritized $a$ for determining $a^{c}$, and the scores were always high. In other words, this behavior suggests a lack of exploration, although this will be evident later when we look at the behavior of the learned policies.
@@ IV-D Results
To compare the five methods more quantitatively, two criteria are statistically evaluated. The one shows the ability to maintain the control performance during data collection, labeled Retention, which is computed as the mean of scores during data collection. The other shows the robustness of the learned policies, labeled Robustness, which is evaluated by deploying them under the disturbed environments, where a uniform random disturbance with a probability of 5 % to each action dimension. Note that the maximum intensity of the disturbance was set to the largest action of each dimension for Pusher and HexaAnt, and to the half of it for HalfCheetah due to its instability.
The normalized results of both criteria are summarized in Fig. 4. As the learning process already suggested, EV1 caused a significant performance degradation, while EV2 almost maintained the expert’s control performance. Although the control performance degraded with each proposed modification, it rarely exceeded the worst case of experts (given as the dashed lines). It would simply be expected that the robustness should be inversely proportional to the retention, and this is the case when looking at EV1 and EV2. However, with the similar retention to EV2, C1 was clearly more robust. More surprisingly, C2, which has the better retention than EV1, achieved the same level of robustness, and C3 achieved even better robustness. This indicates that the proposed CubeDAgger could more efficiently facilitate the exploration to improve the robustness.
@@ V Conclusion
This paper proposed a novel IIL method, named CubeDAgger, an improved version of EnsembleDAgger. The improvements are threefold: i) control the output variance of the ensemble model to make the safety decision work better; ii) design an optimization problem to derive consensus from multiple action candidates; and iii) add colored noise for efficient probabilistic exploration. With these improvements, CubeDAgger successfully improved the robustness of learned policies while maintaining dynamic stability during data collection, compared to EnsembleDAgger. Future work will demonstrate the practicality of CubeDAgger through real-world robot experiments and verify that the colored noise is less disruptive to human demonstrations than ordinary white noise.
ACKNOWLEDGMENT
This research was supported by “Strategic Research Projects” grant from ROIS (Research Organization of Information and Systems).

### v2

@@ IV Experiments
IV-A Simulations
In order to statistically verify the performance of the proposed method, CubeDAgger, the following simulations are conducted. The important criteria are threefold: i) the difference between the expert’s action $a$ and the executed action $a^{c}$ during the data collection process; ii) the degree of degradation of control performance during the data collection process; iii) the robustness of the trained policies.
IV-A1 Tasks
Three dynamic tasks with different robots simulated in Mujoco are performed. The first Pusher, where a manipulator pushes an object to its destination without grasping it, might cause the object to slide into hard-to-push positions by wrong actions. The second HalfCheetah, where a 2D legged robot moves forward while standing on its head [^2] This behavior was achieved by changing the weights of rewards, is unstable and sensitive to mistakes. The third HexaAnt, which has a multi-degree-of-freedom system with three joints in each of its six legs for walking foward, is easy to intervene in state transitions and stability with various action patterns. These tasks were trained multiple times with the RL algorithm [17], and the best policies are selected as their experts. Since the experts maximized the sum of rewards, the control performance is evaluated by it.
IV-A2 Models
The policy of the imitation agent is approximated by neural networks with two fully-connected layers (each with 100 neurons) in the hidden layer. To make it a lightweight ensemble model, only the output layer is separated to take $K=10$ outputs (i.e. the mean and log scale of Gaussian distribution), while sharing the input and hidden layers [27]. This policy ensembles (and other learnable parameters) are optimized using AdaTerm [14], a stochastic gradient descent method that is robust to noise and outliers, with its default configuration. The dataset $D$ is assumed to have the sufficient capacity to hold all the data during the experiments, and all the data in $D$ are uniformly randomly replayed once at the end of each episode with batch size $B=50$ each.
IV-A3 Comparisons
@@
IV-A4 Results
The difference $|a-a^{c}|$ (labeled Diff) and the sum of rewards (labeled Score) at the time of data collection are plotted in Fig. 4. Apparently, EV1 frequently performed actions different from those of the experts, causing large drops in the scores. This behavior would involve significant risk during data collection in real-world robot applications, even if it may be needed to learn how to recover from (accumulated) errors in case of failure. On the other hand, EV2 mostly prioritized $a$ for determining $a^{c}$, and the scores were always high. In other words, this behavior suggests a lack of exploration, although this will be evident later when we look at the behavior of the trained policies.
@@ IV-A4 Results
To compare the five methods more quantitatively, two criteria are statistically evaluated. The one shows the ability to maintain the control performance during data collection, labeled Retention, which is computed as the mean of scores during data collection. The other shows the robustness of the trained policies, labeled Robustness, which is evaluated by deploying them under the disturbed environments, where a uniform random disturbance with a probability of 5 % to each action dimension. Note that the maximum intensity of the disturbance was set to the largest action of each dimension for Pusher and HexaAnt, and to the half of it for HalfCheetah due to its instability.
The normalized results of both criteria are summarized in Fig. 5. As the learning process already suggested, EV1 caused a significant performance degradation, while EV2 almost maintained the expert’s control performance. Although the control performance degraded with each proposed modification, it rarely exceeded the worst case of experts (given as the dashed lines). It would simply be expected that the robustness should be inversely proportional to the retention, and this is the case when looking at EV1 and EV2. However, with the similar retention to EV2, C1 was clearly more robust. More surprisingly, C2, which has the better retention than EV1, achieved the same level of robustness, and C3 achieved even better robustness. This indicates that the proposed CubeDAgger could more efficiently facilitate the exploration to improve the robustness without violating the dynamic stability.
@@
IV-B Scooping balls with quadruped robot
To demonstrate the practicality of CubeDAgger, real-world experiments are conducted. Since a human serves as the expert, the following three points are confirmed: i) EnsembleDAgger’s switching system hinders proper human supervision; ii) Exploration with time-inconsistent white noise impairs human operability; iii) CubeDAgger mitigates these issues and enables stable imitation from human.
IV-B1 Conditions
The experiments use a quadruped robot developed by Ahead.IO (see Fig. 1). Its torso pose is kinematically controlled, so the action space is defined as six dimensions for its target. Note that the action is sent to the robot 30 Hz and smoothed in the internal controller, which runs 90 Hz. A net is attached to the lower of the torso, with upper and lower boundaries of the action space established to ensure the net reached the ground and the entire workspace. A Intel RealSense D435i is mounted at the front of the torso to capture 640x480 RGB images at 30 Hz. The human expert views them and operate the robot via a SpaceMouse.
At the start of each episode, three balls are positioned around the net within the range captured. They are scooped in the order of front, right, and left, although this order is not necessarily satisfied due to disturbances caused by the agent’s actions. Each episode is set to one minute (i.e. 1800 steps). Learning is completed over 30 episodes (i.e. 30 minutes), with the agent’s policy being evaluated every five episodes.
To accomplish this task, its state space is defined as follows. First, the six-dimensional target pose in the internal controller is required to predict the next target pose according to the given action. Second, the RGB image is reshaped to 518x518 and converted to the corresponding 384-dimensional visual features using DINOv2 [26] (more specifically, vit_small_patch14_dinov2.lvd142m in Hugging Face). Finally, to perceive the dynamical features, the above two are fed into an echo state network [15] with 512 neurons. In total, the state space has 902 dimensions.
The model and learning conditions for the agent’s policy are mostly the same as those in the simulations, but the following three points differ. To avoid underfitting with the larger state dimension, the neurons of the two hidden layers are increased to 512 and 256, respectively. To carefully adapt to more diverse situations, the learning rate of AdaTerm is reduced $1/10$ times (i.e. $10^{-4}$). To achieve stable real-time control at 30 Hz, the number of optimization iterations of ITP method is reduced by half (i.e. eight).
IV-B2 Results
The experiments are summarized in the attached video. First, with EnsembleDAgger with $\bar{\Delta}=1$ and $\bar{\sigma}=0.1$, the agent monopolized control authority early in episodes, making erroneous motions. To gain the control authority, the human expert intentionally gave excessive actions, but the actions suddenly switched caused such as collisions with walls due to the delay in handling them.
CubeDAgger with white noise for exploration (i.e. $T\to 0$) could eliminate the dangerous switching behaviors. However, the oscillatory behaviors exhibited by the agent made it difficult for the human expert to provide optimal operation. Furthermore, such oscillations affected the observed images; the screen, constantly fluctuating minutely, caused motion sickness in the human expert.
In contrast, the proposed CubeDAgger successfully executed all episodes safely while promoting diversity in situations through moderate perturbations. While the above two baselines were truncated only with one trial for each due to heavy burden on the human expert, the proposed method could be conducted three trials comfortably. Indeed, only two balls could not be scooped in total episodes.
More quantitatively, the test results of the agent policies conducted every five episodes are depicted in Fig. 6. Consistent with subjective ease of operation, the average $|a-a^{c}|$ every 5 minutes was smallest with the proposed method. In addition, the baselines never scoop all three balls even once, whereas the proposed method eventually succeeded in scooping all balls in all three trials.
Furthermore, when the acquired policies by the proposed method were tested five times, the success rate of scooping the three balls was $8/15$. Remarkably, Among the successful eight episodes, some performed error recovery motions, indicating the benefits of IIL. In the failed seven episodes, while two balls could be scooped, most cases involved misrecogniting the state as allowing three balls to be scooped: namely, the last ball was occluded behind the net in the images. Alternatively, the unusual orders mixed in with the supervision data were executed, leading to the situations where the policies were insufficiently optimized.
Figure 6: Evaluation of real-world scooping task
@@ V Conclusion
This paper proposed a novel IIL method, named CubeDAgger, an improved version of EnsembleDAgger, for making it applicable to dynamic tasks. The improvements are threefold: i) control the output variance of the ensemble model to make the safety decision work better; ii) design an optimization problem to derive consensus from multiple action candidates; and iii) add colored noise for efficient stochastic exploration. With these improvements, CubeDAgger successfully improved the robustness of trained policies while maintaining dynamic stability during data collection, compared to EnsembleDAgger. In the real-world scooping demonstration, CubeDAgger successfully acquired robust policies while safely collecting data without excessively interfering with the human operator.
However, as found in the experiments, human experts in dynamic systems cannot always select optimal actions, leading to noisy data. To mitigate this adverse effect, it should be desirable to integrate noise-robust learning algorithms and/or data selection methods.